OpenStudy (anonymous):

Which statement is true about the sum square root of 50 + square root of 6? It is rational and equal to 2. It is rational and equal to 8. It is irrational and equal to 2square root of 2. It is irrational and equal to 8square root of 2.

OpenStudy (anonymous):

@binks

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

\[\sqrt{50}+\sqrt{6}=5\sqrt{2}+\sqrt{6}\] I can't find it matching to any of the above statements. The above surd cannot be simplified any further

OpenStudy (anonymous):

actually the 6 is supposed to be 18 im not sure why it came up as 6 so sorry

OpenStudy (anonymous):

haha okay that makes more sense okay so we have:\[\sqrt{50}+\sqrt{18}=\sqrt{25*2}+\sqrt{9*2}=5\sqrt{2}+3\sqrt{2}\] just add the coefficients and if it cannot be expressed as a fraction then it is irrational [hint, a surd is not a fraction]! Hope that helps!

OpenStudy (anonymous):

how do you add all of it?

OpenStudy (anonymous):

You just add the coefficients :-) so:\[5\sqrt{2}+3\sqrt{2} = (5+3)\sqrt{2}= ?\]

OpenStudy (anonymous):

so its D right

OpenStudy (anonymous):

yup